OpenStudy (anonymous):

Given the problem (-3)^2 x (-3)^4, which is the correct answer, (-3)^6 or (9)^6? Explain how you got your answer.

OpenStudy (ghazi):

(x)^2*(x)^2= (x)^2+2=(x)^4...so (-3)^6

OpenStudy (cwrw238):

yup

OpenStudy (ghazi):

\[X^a*X^b=X^{a+b}\]